OpenStudy (anonymous):

The function f(x) = 582(7)x represents the growth of a mosquito population every year in a remote swamp. Troy wants to manipulate the formula to an equivalent form that calculates every 2 months, not every year. Which function is correct for Troy's purposes?

OpenStudy (anonymous):

f(x) = 582(7)^x

OpenStudy (the_fizicx99):

plug in for x then multiply it by 7 then 582

OpenStudy (the_fizicx99):

Whatever you get for 7*^x multiply it by 582

OpenStudy (anonymous):

28518?

OpenStudy (the_fizicx99):

Could you show your process?

OpenStudy (anonymous):

im so confused :/ these are my options f(x) = 582(7 1/6)12x f(x) = 52(7)^x f(x) = 582(7^12)^x/12 f(x) = 5852(7)^x

OpenStudy (anonymous):

is it the first one

OpenStudy (the_fizicx99):

Says 2 years, but calculate them in 2 months, so it would be ^12 1/6 = 2 months

OpenStudy (the_fizicx99):

I'd say the first one.

OpenStudy (anonymous):

i agree :)

OpenStudy (anonymous):

i have more problems i need help with, could you help?

OpenStudy (the_fizicx99):

I have to go eat, sorry :/

OpenStudy (anonymous):

ok thanks anyways

OpenStudy (the_fizicx99):

Np :)